A 70kg box is slid along a horizontal floor by a horizontal 400Nt force. Find the distance travelled by the box in 5 seconds if

a) there is no friction

b) coefficient of friction = 0.20


1
Expert's answer
2022-04-11T12:03:16-0400

(a)

"a=\frac{F_{net}}{m}=\frac{400}{70}=5.7\:\rm m/s^2"

"d=at^2/2=5.7*5^2/2=71\;\rm m"

(b)

"a=\frac{F_{net}}{m}=\frac{400-70*9.8*0.2}{70}=3.75\:\rm m/s^2"

"d=at^2/2=3.75*5^2/2=47\;\rm m"